Consider two vectorsб vector a and vector bб such that |vector a|=|vector b|=2.Their scalar product vector a . Vector b is equal to 2 as well. What is the angle between vector a and vector b,
1
Expert's answer
2018-05-27T08:37:08-0400
We use the formula for the angle between vector a and vector b:
cos⁡α=(a∙b)/ab=2/(2∙2)=1/2
α=cos^(-1)⁡〖1/2〗=60°.